# Bicategories of C*-correspondences as Dwyer–Kan localisations

### v1

@@ 2. The nerve of the proper correspondence bicategory
    (2.1) $\mathcal{E}_{ij}\otimes_{A_{j}}(\mathcal{E}_{jk}\otimes_{A_{k}}\mathcal{E}_{kl})$$(\mathcal{E}_{ij}\otimes_{A_{j}}\mathcal{E}_{jk})\otimes_{A_{k}}\mathcal{E}_{kl}$$\mathcal{E}_{ij}\otimes_{A_{j}}\mathcal{E}_{jl}$$\mathcal{E}_{il}$$\mathcal{E}_{ik}\otimes_{A_{k}}\mathcal{E}_{kl}$$\mathrm{ass}$$\mathrm{id}_{\mathcal{E}_{ij}}\otimes_{A_{j}}u_{jkl}$$u_{ijk}\otimes_{A_{k}}\mathrm{id}_{\mathcal{E}_{kl}}$$u_{ijl}$$u_{ikl}$
@@
(2.2) $\Gamma_{\varphi}\otimes_{B}\Gamma_{\psi}\cong\Gamma_{\psi\circ\varphi},\qquad b\otimes c\mapsto\psi(b)c;$
this map is isometric, and $\psi\circ\varphi(A)\psi(B)C=\psi\circ\varphi(A)C$ implies that it is surjective. The isomorphisms in (2.2) make $\varphi\mapsto\Gamma_{\varphi}$ into a homomorphism of bicategories $\mathcal{C}^{*}_{+}\to\mathfrak{Corr}_{\mathrm{prop}}$. So it induces a simplicial map $\Gamma\colon N\mathcal{C}^{*}_{+}\to N\mathfrak{Corr}_{\mathrm{prop}}$.
@@ 3. $\mathrm{C}^{*}$-stable functors
into a Hilbert module over $\mathbb{K}(\mathcal{E}\oplus B)$ that contains $\mathbb{K}(\mathcal{E}\oplus B)$ as a direct summand, and the compact operators on this Hilbert module are $\mathbb{K}(\mathcal{E}\oplus B\oplus\ell^{2}(\mathbb{N}_{\geq 1},B))$. Thus $i_{2}$ is also a corner embedding. The argument above provides another extension-by-zero map $i_{4}\colon\mathbb{K}(\mathcal{E}\oplus B\oplus\ell^{2}(\mathbb{N}_{\geq 1},B))\to\mathbb{K}((\mathcal{E}\oplus B)\otimes\ell^{2}(\mathbb{N}))$ so that $i_{5}\mathrel{\vcentcolon=}i_{4}\circ i_{2}$ is a special corner embedding, and then $F(i_{4})\circ F(i_{2})$ is an equivalence. Kasparov’s Stabilisation Theorem implies $(\mathcal{E}\oplus B)\otimes\ell^{2}(\mathbb{N})\cong B\otimes\ell^{2}(\mathbb{N})$, and so the composite $i_{4}\circ i_{2}\circ i_{1}$ is a special corner embedding as well. Therefore, $F(i_{4})\circ F(i_{2})\circ F(i_{1})$ is an equivalence as well. Now the Two-Out-Of-Six Lemma implies that $F(i_{1})$ is an equivalence. ∎
@@ 4. Proof of the main theorem
(4.1) $\mathcal{E}_{S}\mathrel{\vcentcolon=}\mathcal{E}_{i_{0},i_{k}}\oplus\mathcal{E}_{i_{1},i_{k}}\oplus\dotsb\oplus\mathcal{E}_{i_{k},i_{k}},$
@@
extended by zero. These operators are equal by (2.1). ∎
@@
in $\mathcal{D}$. This gives a recipe to construct $\bar{F}(\mathcal{E})$. If $F$ is $\mathrm{C}^{*}$-stable, then $F(i_{\mathcal{E}})$ is an equivalence in $\mathcal{D}$. We may use a quasi-inverse of it to complete the pair of maps $F(f_{\mathcal{E}})$, $F(i_{\mathcal{E}})$ to a triangle as in (4.2). Hence we get $\bar{F}(\mathcal{E})$ for all arrows in $\mathfrak{Corr}_{\mathrm{prop}}$, together with triangles $\bar{F}(u_{\mathcal{E}})$.
@@
(4.3) $\bar{G}(A_{\varphi(i)},\mathcal{E}_{\varphi(i)\varphi(j)},u_{\varphi(i)\varphi(j)\varphi(k)})=\bar{G}(A_{i},\mathcal{E}_{ij},u_{ijk})\circ\varphi_{*}.$
If the given simplex $(A_{i},\mathcal{E}_{ij},u_{ijk})$ is degenerate, then (4.3) for surjective maps determines $\bar{G}$ by its values on simplices in $N\mathfrak{Corr}_{\mathrm{prop}}$ of smaller dimension. Since we are going to construct $\bar{G}$ inductively over the dimension of the simplices, there is nothing to do for degenerate simplices and we may assume from now on that $(A_{i},\mathcal{E}_{ij},u_{ijk})$ is nondegenerate. In addition, it then suffices to check (4.3) for injective maps.
The condition (4.3) determines
(4.4) $\bar{G}(i_{0},\dotsc,i_{k},S_{k+1},\dotsc,S_{\ell})$
if $S_{\ell}\neq[n]$ because then there is $\varphi\colon[n-1]\to[n]$ with $(i_{0},\dotsc,S_{\ell})\in\varphi_{*}(\operatorname{\overline{Bar}}{\vartriangle^{\!n-1}})$. Since we want $\bar{G}$ to extend $G$, the condition (4.3) determines (4.4) if $k\leq 0$. These simplices form a subcomplex of $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ on which $\bar{G}$ is already given. Notice that the prescription for $\bar{G}$ on this subcomplex is a simplicial map. We must extend it to $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$.
It remains to define (4.4) for simplices in $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ with $k\geq 1$ and $S_{\ell}=[n]$. We do this by a recursion through $k=1,\dotsc,n$. When constructing (4.4) for $k\geq 1$, we assume that $\bar{G}$ is already given on simplices in $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ with smaller value of $k$ (which is indeed so for $k=1$). For fixed $k$, we run another recursion on $\ell$. In each step, we fill a horn to construct (4.4) if $S_{k+1}=\{i_{0},\dotsc,i_{k}\}$ and $S_{k+1}\subsetneq S_{k+2}\subsetneq\cdots\subsetneq S_{\ell}=[n]$, assuming that (4.4) has been constructed for smaller values of $\ell$.
@@
(4.5) $(i_{0},\dotsc,i_{k},S_{k+1},\dotsc,S_{\ell})$
for $0\leq j\leq k$ has a smaller value of $k$, so that $\bar{G}$ on this face has been constructed. The $j$th face of (4.5) for $k+2\leq j\leq\ell$ has a smaller value of $\ell$ and still contains $S_{k+1}=\{i_{0},\dotsc,i_{k}\}$, so that $\bar{G}$ on this face has been constructed. The values of $\bar{G}$ on these boundary faces give an $(\ell,k+1)$-horn in $\mathcal{D}$. If $\ell>k+1$, this is an inner horn and so it may be filled in the quasi-category $\mathcal{D}$. Let $\ell=k+1$. The arrow $i_{k}\to S_{k+1}=\{i_{0},\dotsc,i_{k}\}$ induces a corner embedding, and $G$ maps this to an equivalence in $\mathcal{D}$ by assumption; hence we have got a special $(\ell,\ell)$-horn, and these may be filled in all quasi-categories (see [Joyal:Quasi-categories]*Theorem 1.3). (By the way, the case $\ell=k+1$ only occurs if $k=n$ because $S_{\ell}=[n]$.)
The above recursion for fixed $k$ gives (4.4) for all nondegenerate simplices in $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ with $S_{k+1}=\{i_{0},\dotsc,i_{k}\}$ and $S_{\ell}=[n]$. It also gives $\bar{G}$ on the missing k+1k+1st face
@@
Hence our recursion defines $\bar{G}(i_{0},\dotsc,S_{\ell})$ for all nondegenerate simplices in $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ with $S_{\ell}=[n]$. We extend the map to degenerate simplices in the unique way dictated by the degeneracy maps and take this together with the values already known for $S_{\ell}\subsetneq[n]$. This defines $\bar{G}$ on all simplices of $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$. Since $\bar{G}$ satisfies (4.3),
@@
For $n=1$, there is only one recursion step $k=1$, $\ell=2$, which requires one $(2,2)$-horn to be filled. This is exactly the construction of the triangle in (4.2). The value $\bar{F}(\mathcal{E})$ on the edge $\mathcal{E}$ in $N\mathfrak{Corr}_{\mathrm{prop}}$ is indicated in (4.2).

### v2

@@ 2. The nerve of the proper correspondence bicategory
    (1) $\mathcal{E}_{ij}\otimes_{A_{j}}(\mathcal{E}_{jk}\otimes_{A_{k}}\mathcal{E}_{kl})$$(\mathcal{E}_{ij}\otimes_{A_{j}}\mathcal{E}_{jk})\otimes_{A_{k}}\mathcal{E}_{kl}$$\mathcal{E}_{ij}\otimes_{A_{j}}\mathcal{E}_{jl}$$\mathcal{E}_{il}$$\mathcal{E}_{ik}\otimes_{A_{k}}\mathcal{E}_{kl}$$\mathrm{ass}$$\mathrm{id}_{\mathcal{E}_{ij}}\otimes_{A_{j}}u_{jkl}$$u_{ijk}\otimes_{A_{k}}\mathrm{id}_{\mathcal{E}_{kl}}$$u_{ijl}$$u_{ikl}$
@@
(2) $\Gamma_{\varphi}\otimes_{B}\Gamma_{\psi}\cong\Gamma_{\psi\circ\varphi},\qquad b\otimes c\mapsto\psi(b)c;$
this map is isometric, and $\psi\circ\varphi(A)\psi(B)C=\psi\circ\varphi(A)C$ implies that it is surjective. The isomorphisms in (2) make $\varphi\mapsto\Gamma_{\varphi}$ into a homomorphism of bicategories $\mathcal{C}^{*}_{+}\to\mathfrak{Corr}_{\mathrm{prop}}$. So it induces a simplicial map $\Gamma\colon N\mathcal{C}^{*}_{+}\to N\mathfrak{Corr}_{\mathrm{prop}}$.
@@ 3. $\mathrm{C}^{*}$-stable functors
into a Hilbert module over $\mathbb{K}(\mathcal{E}\oplus B)$ that contains $\mathbb{K}(\mathcal{E}\oplus B)$ as a direct summand, and the compact operators on this Hilbert module are $\mathbb{K}(\mathcal{E}\oplus B\oplus\ell^{2}(\mathbb{N}_{\geq 1},B))$. Thus $i_{2}$ is also a corner embedding. The argument above provides another extension-by-zero map $i_{4}\colon\mathbb{K}(\mathcal{E}\oplus B\oplus\ell^{2}(\mathbb{N}_{\geq 1},B))\to\mathbb{K}((\mathcal{E}\oplus B)\otimes\ell^{2}(\mathbb{N}))$ so that $i_{5}\mathrel{\vcentcolon=}i_{4}\circ i_{2}$ is a special corner embedding, and then $F(i_{4})\circ F(i_{2})$ is an equivalence. Now the Two-Out-Of-Six Lemma for the composition of $F(i_{4})$, $F(i_{2})$ and $F(i_{1})$ implies that $F(i_{1})$ is an equivalence. ∎
@@ 4. Proof of the main theorem
(3) $\mathcal{E}_{S}\mathrel{\vcentcolon=}\mathcal{E}_{i_{0},i_{k}}\oplus\mathcal{E}_{i_{1},i_{k}}\oplus\dotsb\oplus\mathcal{E}_{i_{k},i_{k}},$
@@
extended by zero. These operators are equal by (1). ∎
@@
in $\mathcal{D}$. This gives a recipe to construct $\bar{F}(\mathcal{E})$. If $F$ is $\mathrm{C}^{*}$-stable, then $F(i_{\mathcal{E}})$ is an equivalence in $\mathcal{D}$. We may use a quasi-inverse of it to complete the pair of maps $F(f_{\mathcal{E}})$, $F(i_{\mathcal{E}})$ to a triangle as in (4). Hence we get $\bar{F}(\mathcal{E})$ for all arrows in $\mathfrak{Corr}_{\mathrm{prop}}$, together with triangles $\bar{F}(u_{\mathcal{E}})$.
@@
(5) $\bar{G}(A_{\varphi(i)},\mathcal{E}_{\varphi(i)\varphi(j)},u_{\varphi(i)\varphi(j)\varphi(k)})=\bar{G}(A_{i},\mathcal{E}_{ij},u_{ijk})\circ\varphi_{*}.$
If the given simplex $(A_{i},\mathcal{E}_{ij},u_{ijk})$ is degenerate, then (5) for surjective maps $\varphi$ determines $\bar{G}$ by its values on simplices in $N\mathfrak{Corr}_{\mathrm{prop}}$ of smaller dimension. Since we are going to construct $\bar{G}$ inductively over the dimension of the simplices, there is nothing to do for degenerate simplices and we may assume from now on that $(A_{i},\mathcal{E}_{ij},u_{ijk})$ is nondegenerate. In addition, it then suffices to check (5) for injective maps.
The condition (5) determines
(6) $\bar{G}(i_{0},\dotsc,i_{k},S_{k+1},\dotsc,S_{\ell})$
if $S_{\ell}\neq[n]$ because then there is $\varphi\colon[n-1]\to[n]$ with $(i_{0},\dotsc,S_{\ell})\in\varphi_{*}(\operatorname{\overline{Bar}}{\vartriangle^{\!n-1}})$. Since $\bar{G}$ must extend $G$, we already know (6) if $k\leq 0$. These simplices form a subcomplex of $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$. The prescription for $\bar{G}$ on this subcomplex is a simplicial map. We must extend it to $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$.
It remains to define (6) for simplices in $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ with $k\geq 1$ and $S_{\ell}=[n]$. We do this by a recursion through $k=1,\dotsc,n$. When constructing (6) for $k\geq 1$, we assume that $\bar{G}$ is already given on simplices in $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ with smaller value of $k$ (which is indeed so for $k=1$). For fixed $k$, we run another recursion on $\ell$. In each step, we fill a horn to construct (6) if $S_{k+1}=\{i_{0},\dotsc,i_{k}\}$ and $S_{k+1}\subsetneq S_{k+2}\subsetneq\cdots\subsetneq S_{\ell}=[n]$, assuming that (6) has been constructed for smaller values of $\ell$.
@@
(7) $(i_{0},\dotsc,i_{k},S_{k+1},\dotsc,S_{\ell})$
for $0\leq j\leq k$ has a smaller value of $k$, so that $\bar{G}$ on this face has been constructed. The $j$th face of (7) for $k+2\leq j\leq\ell$ has a smaller value of $\ell$ and still contains $S_{k+1}=\{i_{0},\dotsc,i_{k}\}$, so that $\bar{G}$ on this face has been constructed. The values of $\bar{G}$ on these boundary faces give an $(\ell,k+1)$-horn in $\mathcal{D}$. If $\ell>k+1$, this is an inner horn and so it may be filled in the quasi-category $\mathcal{D}$. Let $\ell=k+1$. The arrow $i_{k}\to S_{k+1}=\{i_{0},\dotsc,i_{k}\}$ induces a corner embedding, and $G$ maps this to an equivalence in $\mathcal{D}$ by assumption; hence we have got a special $(\ell,\ell)$-horn, and these may be filled in all quasi-categories (see [Joyal:Quasi-categories]*Theorem 1.3). (By the way, the case $\ell=k+1$ only occurs if $k=n$ because $S_{\ell}=[n]$.)
The above recursion for fixed $k$ gives (6) for all nondegenerate simplices in $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ with $S_{k+1}=\{i_{0},\dotsc,i_{k}\}$ and $S_{\ell}=[n]$. It also gives $\bar{G}$ on the missing k+1k+1st face
@@
Hence our recursion defines $\bar{G}(i_{0},\dotsc,S_{\ell})$ for all nondegenerate simplices in $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$ with $S_{\ell}=[n]$. We extend the map to degenerate simplices in the unique way dictated by the degeneracy maps and take this together with the values already known for $S_{\ell}\subsetneq[n]$. This defines $\bar{G}$ on all simplices of $\operatorname{\overline{Bar}}{\vartriangle^{\!n}}$. Since $\bar{G}$ satisfies (5),
@@
For $n=1$, there is only one recursion step $k=1$, $\ell=2$, which requires one $(2,2)$-horn to be filled. This is exactly the construction of the triangle in (4). The value $\bar{F}(\mathcal{E})$ on the edge $\mathcal{E}$ in $N\mathfrak{Corr}_{\mathrm{prop}}$ is indicated in (4).